Given an A.P. whose terms are all positive integers. The sum of its first nine terms is greater then 200 and less than 220. If the second term in it is 12, then its 4th term is. :

Grade:12

3 Answers

Sher Mohammad IIT Delhi
10 years ago
a+d=12
sum of first 9 terms 200<a+8d<220
200-12<a+8d-a-d<220-12
188<7d<208
27<d <29
taking d=28, 4th term is a+3d=-16+3*28 ,

sher mohammad
b.tech, iit delhi





sai krishna
10 years ago
sum of n terms in an A.P = N/2[2a+(n-1)d] So, sum to 9 terms= 9/2[2a+8d] According to given question, 200<9/2[2a+8d]<220 = 200<9[a+4d]<220 = 200/9
sai krishna
10 years ago
sum of n terms in an A.P = N/2[2a+(n-1)d] So, sum to 9 terms= 9/2[2a+8d] According to given question, 200<9/2[2a+8d]<220 = 200<9[a+4d]<220 = 200/9